# pyMSER - An open-source library for automatic equilibration detection in molecular simulations

### v1

@@ 1 Introduction
Currently, the prevailing practice in the literature is to define a fixed number of equilibration steps, often ranging from 1000 to 50000 [8, 9, 10, 11, 12, 13, 14]. Although for low pressures a smaller number of equilibration steps may be sufficient, one cannot know whether these arbitrarily chosen numbers of steps are sufficient to ensure accurate ensemble averages across all simulation conditions. This becomes especially critical for simulations where the number of molecules within the pore is very large, such as at higher pressures and/or lower temperatures. While some approaches have been proposed to determine the appropriate truncation point in molecular simulations, no single method has yet emerged as the widely accepted or standardized approach.
@@ 2.1 GCMC Simulations
The Grand Canonical Monte Carlo (GCMC) simulations were performed using RASPA (version v2.0.45).[6, 18] Interaction energies between non-bonded atoms were computed via a combination of Lennard-Jones (LJ) and Coulomb potentials
$U_{ij}(r_{ij})=4\varepsilon_{ij}\left[\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}-\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\right]+\frac{1}{4\pi\epsilon}\frac{q_{i}q_{j}}{r_{ij}}$ (1)
where $i$ and $j$ are the atom indexes and $r_{ij}$ is their inter-atomic distance. $\varepsilon_{ij}$ and $\sigma_{ij}$ are the LJ potential wheel depth and zero-energy radius, respectively, and $q_{i}$ and $q_{j}$ are the partial charges of atoms $i$ and $j$. LJ parameters for framework atoms were taken from the Universal Force Field (UFF) [19]. The partial charges were calculated using the EQeq method [20]. The parameters for the adsorbed molecules were taken from the TraPPE [21] force field. The LJ parameters between atoms of different types were calculated using the Lorentz-Berthelot mixing rules.
@@
All simulations were performed with Monte Carlo step sampling such that swap (insertion or deletion with with a probability of 50% each), translations, rotations, and re-insertions were performed with probabilities of 0.5, 0.3, 0.1, and 0.1 respectively. The results that have been presented are analyzed with respect to the number of MC (Monte Carlo) steps. The MC steps were converted from RASPA cycles. Each RASPA cycle contains $max(20,N)$ MC cycles where $N$ is the number of atoms present in the simulation box.
@@ 2.2 The Marginal Standard Error Rule
There are two MSER variants, in addition to the standard version, designed to present better performance on specific use-cases. The first variant, MSER-LLM, aims to preserve as much of the original data series as possible. This is achieved by selecting the Left-most Local Minimum of the MSER curve as the optimal truncation point instead of the global minimum, as in the standard MSER version. This strategy generates a good balance between sample size increase and bias elimination, removing less data than the original method with a minimal impact on the final average accuracy, as will be shown later.
@@ 2.3 Statistical uncertainty
The same process can be carried out for the Standard Error of the mean (SE), defined as $SD/\sqrt{N_{\text{prod}}}$, thus generating the uncorrelated Standard Error of the mean (uSE). Figure 3 illustrates the process of creating uncorrelated batches and its impact on the uncertainty metrics.
@@ 3 Results and discussion
Figure 5: Truncation point detection with different approaches. Both pyMSER (a) and the method suggested by Gowers et al. (b) identify the truncation point around the same value ($\sim 26\cdot 10^{6}$ steps) while the RCA (c) and pyMBAR (d) detects a very late truncation point for the production stage.
@@ 10 ADF_test=True,
Another potential use of pyMSER is to be integrated into simulation workflows, where the simulation is monitored at regular intervals to detect the start of the production stage. Then, a given number of production steps can be performed in order to obtain a well-equilibrated property.
Figure 7: Comparison of approaches for simulating CO2 uptake on MOFs from the CoRE-MOF-2019 database. a) Comparison between the CO2 uptake obtained with two differente approaches. The on-the-fly MSER approach achieves the same results as the standard approach does, albeit with a tiny fraction of the computational cost due to running drastically less GCMC steps. b-c) Histograms comparing execution time using the standard and MSER approach, demonstrating the significant reduction in total execution time achievable with the latter.
To evaluate the potential impact of pyMSER compared to a standard GCMC approach in high-throughput simulations, let us consider the following scenario. A researcher wants to evaluate the carbon capture performance of MOF structures selected from the CoRE-MOF-2019 database [35] by simulating the adsorption of CO2 at 1 bar and 298 K. The researcher has two options: (i) use the standard method from the literature, which involves running a fixed number of MC equilibration steps followed by a fixed number of MC production steps, or (ii) develop a straightforward code that assesses, at specified intervals (e.g., every 10000 steps), whether the simulation has attained equilibrium, and then runs a fixed number of MC production steps.
Figure 7(a) illustrates the uptakes obtained using these two approaches: (i) the standard approach represented on the x-axis, and (ii) the pyMSER approach depicted on the y-axis. Additionally, Figure 7(b) present histograms comparing the execution times of these approaches, highlighting the substantial reduction achievable with the pyMSER method. On average, the execution time using pyMSER is 10 times faster than using the conventional approach, and it can be up to 30 times faster in some cases. In this analysis, we discarded CoRE-MOF-2019 materials whose GCMC simulation took more than 24 hours. The results demonstrate agreement within the standard deviation, indicating that despite its lower computational cost, the pyMSER approach yields consistent results comparable to well-established, albeit more computationally intensive, methodologies.
@@
In general, the uptake and total energy of a single-component gas adsorption simulation reach equilibrium around the same time. Figure S3 illustrates this behavior, showing that both curves reach a plateau around the same period. However, pyMSER analyzes both curves and uses the largest detected $t_{0}$ as the truncation point for selecting the region for adsorption enthalpy calculation.
@@
In this context, pyMSER is an ideal tool for automatically and reliably detecting the optimal truncation point in each simulation, without the need for human intervention or the creation of complex heuristics. pyMSER can also be used as an on-the-fly testing tool in automated workflows, allowing for an objective evaluation of whether a given simulation has reached a certain convergence criterion defined by the user and can be stopped. Possible criteria would be achieving a certain number of data points in the production region or a certain signal-to-noise ratio between the observable value and its uncertainty.
4 Conclusion
@@
Additionally, the MSER statistics do not depend on the nature of the data and are method-agnostic, i.e. independent of the simulation software that provides the data, therefore it can be applied to any time series data, whenever removing initial transient values is necessary.
@@ Acknowledgments
Supporting Information
The Supporting Information is available free of charge on the publisher’s website.

### v2

@@ 1 Introduction
Currently, the prevailing practice in the literature is to define a fixed number of equilibration steps, often ranging from 1,000 to 50,000 [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Although for low pressures a smaller number of equilibration steps may be sufficient, one cannot know whether these arbitrarily chosen numbers of steps are sufficient to ensure accurate ensemble averages across all simulation conditions. This becomes especially critical for simulations where the number of molecules within the pore is very large, such as at higher pressures and/or lower temperatures. While some approaches have been proposed to determine the appropriate truncation point in molecular simulations, no single method has yet emerged as the widely accepted or standardized approach.
@@ 2.1 GCMC Simulations
The Grand Canonical Monte Carlo simulations were performed using RASPA (version v2.0.45).[6, 25] Interaction energies between non-bonded atoms were computed via a combination of Lennard-Jones (LJ) and Coulomb potentials
$U_{ij}(r_{ij})=4\varepsilon_{ij}\left[\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}-\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\right]+\frac{1}{4\pi\epsilon_{0}}\frac{q_{i}q_{j}}{r_{ij}}$ (1)
where $i$ and $j$ are the atom indexes and $r_{ij}$ is their inter-atomic distance. $\varepsilon_{ij}$ and $\sigma_{ij}$ are the LJ potential well depth and zero-energy radius, respectively, $q_{i}$ and $q_{j}$ are the partial charges of atoms $i$ and $j$, and $\epsilon_{0}$ is the electric constant. LJ parameters for framework atoms were taken from the Universal Force Field (UFF) with the values provided in Supporting Information [26]. The partial charges were calculated using the EQeq method [27]. The parameters for the adsorbed molecules were taken from the TraPPE [28] force field. The LJ parameters between atoms of different types were calculated using the Lorentz-Berthelot mixing rules.
@@
All simulations were performed with Monte Carlo step sampling such that swap (insertions or deletions), translations, rotations, and re-insertions were performed with probabilities of 0.5, 0.3, 0.1, and 0.1 respectively. The results that have been presented are analyzed with respect to the number of MC steps. The MC steps were converted from RASPA cycles. Each RASPA cycle contains $max(20,N)$ MC cycles where $N$ is the number of atoms present in the simulation box.
@@ 2.2 The Marginal Standard Error Rule
where $n$ represents the number of points in the original series, $Y_{i}$ denotes the property value at the $i$-th point, and $\bar{Y}_{n,k}$ the average of the remaining data points.
@@
There are two MSER variants, in addition to the original version, designed to present better performance on specific use-cases. The first variant, MSER-LLM, aims to preserve as much of the original data series as possible. This is achieved by selecting the Left-most Local Minimum of the MSER curve as the optimal truncation point instead of the global minimum, as in the original MSER version. This strategy generates a good balance between sample size increase and bias elimination, removing less data than the original method with a minimal impact on the final average accuracy, as will be shown later.
@@ 2.3 Statistical uncertainty
where $\tau$ represents the “autocorrelation time” and $Y(x)$ denotes the property value at the $x$-th point.
@@
The same process can be carried out for the Standard Error of the mean (SE), defined as $SD/\sqrt{N_{\text{prod}}}$, thus generating the uncorrelated Standard Error of the mean (uSE). Figure 3 illustrates the process of creating uncorrelated batches and its impact on the uncertainty metrics. We note that the selection of the uncertainty metric employed in the calculation does not impact the estimation of the truncation point or the average value.
@@ 3 Results and discussion
Figure 5: Truncation point detection with different approaches. Both pyMSER (a) and the method suggested by Gowers et al. (b) identify the truncation point around the same value ($\sim 26\cdot 10^{6}$ steps) while the RCA (c) and pyMBAR (d) detects a very late truncation point for the production stage. These methodologies were applied to GCMC simulation data for H2O adsorption in MOF HKUST-1 at 273K.
@@
An alternative to the ADF test is the Gelman-Rubin method [40], which can also be used to verify the equilibration of the production stage. We included an in-depth analysis of the Gelman-Rubin method to the Supporting Information section.
@@ 10 ADF_test=True,
Another potential use of pyMSER is to be integrated into simulation workflows, where the simulation is monitored at regular intervals to detect the start of the production stage. Then, a given number of production steps can be performed in order to obtain a well-equilibrated property. An example of this on-the-fly analysis workflow is implemented in the GitHub repository provided in the Data and Software Availability section.
Figure 7: Comparison of approaches for simulating CO2 uptake on MOFs from the CoRE-MOF-2019 database. a) Comparison between the CO2 uptake obtained with two differente approaches. The on-the-fly MSER approach achieves the same results as the standard approach does, albeit with a tiny fraction of the computational cost due to running drastically less MC steps. b-c) Histograms comparing execution time using the standard and MSER approaches, demonstrating the significant reduction in total execution time achievable with the latter.
To evaluate the potential impact of pyMSER compared to the standard GCMC approach widely used in high-throughput simulations, let us consider the following scenario. A researcher wants to evaluate the carbon capture performance of MOF structures selected from the CoRE-MOF-2019 database [43] by simulating the adsorption of CO2 at 1 bar and 298 K. The researcher has two options: (i) use the standard method from the literature, which involves running a fixed number of MC equilibration cycles followed by a fixed number of MC production cycles, or (ii) develop a straightforward code that assesses, at specified intervals (e.g., every 1000 cycles), whether the simulation has attained equilibrium, and then runs a fixed number of MC production cycles.
We take, as the “standard” approach, one that finds ample usage in the literature. [15, 16, 17, 18, 19, 20, 21] Figure 7(a) illustrates the uptakes obtained using these two approaches: (i) the standard approach represented on the x-axis, with 50,000 equilibration cycles and 50,000 production cycles, and (ii) the pyMSER approach depicted on the y-axis, with 10,000 production cycles simulated after reaching equilibration. Additionally, Figure 7(b) present histograms comparing the execution times of these approaches, highlighting the substantial reduction achievable with the pyMSER method. On average, the execution time using pyMSER is $\frac{1}{10}$ that of the standard approach, and it can be up to $\frac{1}{30}$ in some cases. In this analysis, we discarded CoRE-MOF-2019 materials whose GCMC simulation took more than 24 hours due to limitations in the job submission queue system. The results demonstrate agreement within the standard deviation, indicating that despite its lower computational cost, the pyMSER approach yields consistent results comparable to well-established, albeit more computationally intensive, methodologies.
@@
In general, the uptake and total energy of a single-component gas adsorption simulation do not necessarily reach equilibrium at exactly the same time. Therefore, pyMSER analyzes both time series independently and uses the largest detected $t_{\text{eq}}$ as the truncation point for selecting the region to calculate the adsorption enthalpy. Figure S3 illustrates this behavior.
@@
In this context, pyMSER is an ideal tool for automatically and reliably detecting the optimal truncation point in each simulation, without the need for human intervention or the creation of complex heuristics. pyMSER can also be used as an on-the-fly testing tool in automated workflows, allowing for an objective evaluation of whether a given simulation has reached a certain convergence criterion defined by the user and can be stopped. Possible criteria would be achieving a certain number of data points in the production region or a certain signal-to-noise ratio between the observable value and its uncertainty. We refer readers to the GitHub repository where they can find an example workflow depicting on-the-fly pyMSER analysis, which can serve as a starting point for the development of more complex workflows.
Additionally, the MSER statistics do not depend on the nature of the data and are method-agnostic, i.e. independent of the simulation software that provides the data, therefore it can be applied to any time series data, whenever removing initial transient values is necessary. While extensive testing demonstrates pyMSER’s robustness, its use beyond the scope explored here requires independent evaluation by the user. For this purpose, we recommend the application of the Augmented Dickey-Fuller test, as implemented in pyMSER, to evaluate the equilibration of the production stage.
4 Conclusions
@@ 29 Cutoff Metrics :
Lennard-Jones parameters
The Lennard-Jones parameters for the UFF force field used in the calculations for the framework atoms are shown in Table S3. The TraPPE and TIP5Ew parameters used for the gas molecules are present in Table S1. The critical parameters used in the Peng-Robinson equation to calculate the fugacity are present in Table S2.
Atom type $\sigma$ (Å) $\varepsilon$ (K) Charge
C_CO2 2.80 27.0 0.700
O_CO2 3.05 79.0 -0.350
N_N2 3.31 36.0 -0.482
N_com - - 0.964
Ow 3.097 89.63 0.964
Hw - - 0.241
Lw - - -0.241
Table S1: Lennard-Jones parameters for molecules.
Gas Critical temperature (K) Critical Pressure (Pa) Acentric factor
CO2 304.1282 7377300.0 0.22394
N2 126.192 3395800.0 0.0372
H2O 304.1282 7377300.0 0.22394
Table S2: Critical parameters for CO2 and N2.
Atom $\varepsilon$ (K) $\sigma$ (Å) Atom $\varepsilon$ (K) $\sigma$ (Å) Atom $\varepsilon$ (K) $\sigma$ (Å)
B_ 47.804 3.582 Ni_ 7.548 2.525 Fr_ 25.16 4.366
C_ 47.854 3.474 Cu_ 2.516 3.114 Ra_ 203.293 3.276
H_ 7.649 2.847 Zn_ 27.676 4.045 La_ 8.554 3.138
N_ 38.948 3.263 In_ 276.76 4.09 Ce_ 6.542 3.169
O_ 48.156 3.034 Sn_ 276.76 3.983 Pr_ 5.032 3.213
P_ 161.024 3.698 Sb_ 276.76 3.876 Nd_ 5.032 3.186
S_ 173.101 3.591 Te_ 286.824 3.769 Pm_ 4.529 3.161
F_ 36.482 3.094 Rb_ 20.128 3.666 Sm_ 4.026 3.137
I_ 256.632 3.698 Sr_ 118.252 3.244 Eu_ 4.026 3.112
K_ 17.612 3.397 Zr_ 34.721 2.784 Gd_ 4.529 3.001
V_ 8.051 2.801 Nb_ 29.689 2.82 Tb_ 3.522 3.075
W_ 33.714 2.735 Mo_ 28.179 2.72 Dy_ 3.522 3.055
Y_ 36.23 2.981 Tc_ 24.154 2.671 Ho_ 3.522 3.038
U_ 11.07 3.025 Ru_ 28.179 2.64 Er_ 3.522 3.022
Li_ 12.58 2.184 Rh_ 26.67 2.61 Tm_ 3.019 3.006
Be_ 42.772 2.446 Pd_ 24.154 2.583 Yb_ 114.73 2.99
Al_ 155.992 3.912 Ag_ 18.115 2.805 Lu_ 20.631 3.243
T_ 155.992 3.805 Cd_ 114.73 2.538 Ac_ 16.606 3.099
Si_ 155.992 3.805 Tl_ 342.176 3.873 Th_ 13.083 3.026
Cl_ 142.557 3.52 Pb_ 333.622 3.829 Pa_ 11.07 3.051
Na_ 251.6 2.801 Bi_ 260.658 3.894 Np_ 9.561 3.051
Mg_ 55.855 2.692 Po_ 163.54 4.196 Pu_ 8.051 3.051
Ga_ 201.28 3.912 At_ 142.909 4.233 Am_ 7.045 3.013
Ge_ 201.28 3.805 Rn_ 124.794 4.246 Cm_ 6.542 2.964
As_ 206.312 3.698 Cs_ 22.644 4.025 Bk_ 6.542 2.975
Se_ 216.376 3.591 Ba_ 183.165 3.3 Cf_ 6.542 2.952
Br_ 186.184 3.52 Hf_ 36.23 2.799 Es_ 6.038 2.94
Ca_ 25.16 3.094 Ta_ 40.759 2.825 Fm_ 6.038 2.928
Sc_ 9.561 2.936 Re_ 33.211 2.632 Md_ 5.535 2.917
Ti_ 8.554 2.829 Os_ 18.618 2.78 No_ 5.535 2.894
Cr_ 7.548 2.694 Ir_ 36.734 2.531
Mn_ 6.542 2.638 Pt_ 40.256 2.454
Fe_ 27.676 4.045 Au_ 19.625 2.934
Co_ 7.045 2.559 Hg_ 193.732 2.41
Table S3: Lennard-Jones parameters for UFF force field.
Alternatives to the Augmented Dickey-Fuller test
The Gelman-Rubin factor (GR) method provides a formula for an $R$ factor, which can be used to determine whether a simulation has reached convergence by comparing it to a user-defined criterion, $R^{*}$. In the version of the GR method presented by Anstine et al., the transient stage is excluded before performing the GR analysis. This exclusion is accomplished by fitting an exponential curve to the time series data and identifying the truncation point where the derivative falls below an arbitrary threshold, set at $10^{-4}$. Both the $R^{*}$ value and the derivative threshold were empirically determined through tests on GCMC data.
The application of the above method to GCMC example data is illustrated in Figure S11. For cases with extremely rapid equilibrium, such as (b) and (d), the truncation point detection aligns with the pyMSER method, and the resulting $R$ value is close to 1, indicating correctly that the production data selected is equilibrated.
In the data shown in Figure S11(a), the truncation point is correctly identified. However, the $R$ value is significantly higher than 1.5. This is unexpected, as both visual inspection and the Augmented Dickey-Fuller test suggest that this region is satisfactorily equilibrated. If we had used the standard criterion of $R^{*}=1.5$, we would have incorrectly classified this region as non-equilibrated.
Figure S11: Gelman-Rubin factor applied to several GCMC data. Black lines represent the total uptake, red lines represent the fitted exponential curve, and the purple dots are the equilibration points detected by the Gelman-Rubin factor. The GCMC data corresponds to a) H2O on IRMOF-1 at 300K, b) CO2 on CuBTC MOF at 273K, c) H2O on CuBTC MOF at 300K, d) CO2 on MgMOF74 at 273K, e) H2O on IRMOF-1 at 300K, and f) H2O on IRMOF-1 at 300K.
Cases (c) and (f) presented incorrect values for the truncation point that were specifically caused by fitting an exponential curve to the time series data. In case (c), the data does not precisely follow an exponential model, causing the fitted curve to poorly represent points near the transient-production transition as shown in Figure S12. For case (f), on the start of the fitted curve the derivative is smaller than the imposed threshold of $10^{-4}$, causing this method to incorrectly select a premature truncation point. Thus, the calculated $R$ value is also incorrect and, consequently, any conclusions drawn from this analysis will also be incorrect.
Figure S12: Detection of the truncation point based on the fit of an exponential curve on the uptake data. The threshold for the derivative curve leads to detecting an incorrect truncation point.
For case (e), the large $R$ value indicates that the curve did not reach equilibrium, and neither is it possible to fit an exponential curve to the data. This results in the detection of the truncation point to be extremely premature.
This, again, reinforces one of the value propositions of pyMSER: a parameter-free method that analyses the whole time series to minimise an error function. Even though the GR method has proven to be useful in many scenarios, the dependence on carefully-determined ad hoc parameters makes it less suitable for an automated high-throughput screening use case. Although extensive tests can be conducted to optimally determine these parameters, one cannot guarantee that they will always work in a scenario were manual data inspection is practically impossible.